You run a regression analysis on a bivariate set of data (n=101n=101). With ¯x=65.2x¯=65.2 and ¯y=42.7y¯=42.7, you obtain the regression equation

y=−1.151x−20.497y=-1.151x-20.497

with a correlation coefficient of r=−0.017r=-0.017. You want to predict what value (on average) for the response variable will be obtained from a value of 190 as the explanatory variable.

What is the predicted response value?
y =
Answer #1

soli 0 Givent biraskite on va The regression analysis set is given bay у = -1.151 x – 20.497; y = -1.1518-20.497 0=-0.017, yo